40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 9 00:34:34 2014, Jenne, Update, LSC, Figuring out where to do DARM->AS55 
    Reply  Tue Sep 9 14:25:46 2014, jamie, Update, LSC, Figuring out where to do DARM->AS55 
    Reply  Wed Sep 10 10:49:39 2014, Manasa, Update, LSC, Y arm green + PSL green mode overlap 
       Reply  Wed Sep 10 21:05:43 2014, Jenne, Update, LSC, Holy sensitivity, Batman! ALS_Y_outOfLoop_10Sept2014.pdfALS_XY_outOfLoop_10Sept2014.pdf
          Reply  Fri Sep 12 00:28:04 2014, ericq, Update, LSC, Holy sensitivity, Batman! ALSoutOfLoop.pdf
Message ID: 10497     Entry time: Fri Sep 12 00:28:04 2014     In reply to: 10491
Author: ericq 
Type: Update 
Category: LSC 
Subject: Holy sensitivity, Batman! 

I took a quick measurement of the ALS stability, using POX and POY as out of loop sensors, using a CARM calibration line to line POX and POY up to the calibrated PHASE_OUT channels at 503Hz. 

  • X arm RMS ~1kHz
    • Could use more low frequency suppression
  • Y arm RMS ~200Hz

ALSoutOfLoop.pdf

ELOG V3.1.3-